OpenStudy (anonymous):

Michael was picking apples and placing them in boxes in a cart. The number of apples that fit in a box is 5 more than the number boxes in a cart. If b represents the number of boxes in each cart, which expression can be used to represent the number of apples Michael picked in 3 carts?

OpenStudy (anonymous):

3b5 3b(b + 5) 3(b + 5) 3b + 5

OpenStudy (anonymous):

i think its B

OpenStudy (anonymous):

x3 - 2y2 - 3x3 + z4 if x = 3, y = 5, and z = -3. first gather and add the same numbers =x3 -3x3 - 2y2 + z4 = -2x3 - 2y2 + z4=-2*3^3 -2*5²+(-3)^4=-54-50+81=-13 2x2 - 2z4 + y2 - x2 + z4 if x = -4, y = 3, and z = 2. =2x2 -x2 - 2z4 + z4 + y2 = x² -z^4 +y² = ( -4)² -2^4 +3²=16 -16 +9 = 9 *****3b(b + 5)*****

OpenStudy (anonymous):

ahh i was right! :D thanks man really appreciate it!

OpenStudy (anonymous):

no problem!